OpenStudy (anonymous):

solve 2x + 6 = 3x +2

OpenStudy (kuoministers):

get it?

OpenStudy (anonymous):

no i dont

OpenStudy (kuoministers):

6 + 2x = 2 + 3x 6 + -1x = 2 + 3x + -3x 6 + -1x = 2 -1x = 2 + -6 x = 4

OpenStudy (anonymous):

ok how about 4h - 6 = 22

OpenStudy (anonymous):

Please post that as a separate question also please close this question and award a medal to whomever helped you most